# Fearless Stochasticity in Expectation Propagation

### v1

@@ Appendix G Computational cost analysis
EP draws $n_{\text{samp}}$ samples from each of the $m$ sites. It then performs $m$ backward mappings, to map from moments back to updated site parameters. The total cost is then $O(mn_{\text{samp}}c_{\text{samp}}+mc_{\text{bwd}})$
@@
EP-$\eta$ draws $n_{\text{samp}}$ samples from each of the $m$ sites. Each update also involves one forward mapping per iteration. In addition, each iteration also requires $m$ JVPs through the backward mapping $\nabla A^{*}(.)$. The primals of this JVP are the same for each site, so the linearisation only needs to be performed once per update, costing $O(2c_{\text{bwd}})$, but we have $m$ cotangents. The overall cost of an EP-$\eta$ iteration is therefore $O(mn_{\text{samp}}c_{\text{samp}}+c_{\text{fwd}}+(2+m)c_{\text{bwd}})$.
@@
EP-$\mu$ draws $n_{\text{samp}}$ samples from each of the $m$ sites. Each update involves one forward mapping, and $m$ backward mappings per iteration, but notably does not require any JVPs. The overall cost of an EP-$\mu$ iteration is therefore $O(mn_{\text{samp}}c_{\text{samp}}+c_{\text{fwd}}+mc_{\text{bwd}})$.
@@
SNEP draws $n_{\text{samp}}$ samples from each of the $m$ sites, and performs $m$ backward mappings per iteration. The total cost is then $O(mn_{\text{samp}}c_{\text{samp}}+mc_{\text{bwd}})$, which is the same as EP.
In this work we largely assume that $c_{\text{samp}}$ is the dominant cost. For example, in the NUTS sampler with default numpyro settings, (as used in our experiments), each sample can involve evaluating up to 1024 gradients of the tilted distribution log density.
The other costs, $c_{\text{fwd}}$ and $c_{\text{bwd}}$, depend on $\mathcal{F}$. In the case of a MVN family with dense covariance, the foward and backward mappings are $O(d^{3})$, for $d$ the dimensionality of $z$. When $d$ is large, $c_{\text{fwd}}$ and $c_{\text{bwd}}$ become more significant, in which case the balance will shift in favour of using more samples per update (and taking larger steps). When $d$ is very large, however, it may be necessary to choose a diagonal covariance base family instead, in which case the mappings can be performed in $O(d)$ time. Also note that when $d$ increases, $c_{\text{samp}}$ will also typically increase at a rate faster than $d$, even with optimal tuning [6]. In the main experiments, we used just a single sample per update for EP-$\eta$ and EP-$\mu$, and they significantly outperformed the baselines when measure in both NUTS steps and wall-clock-time.
We also note that the per-iteration cost of $c_{\text{fwd}}$ and $c_{\text{bwd}}$ could be reduced by exploiting the results of previous iterations. For example, in the case of a MVN with dense covariance, the cost of the forward and backward mappings are dominated by matrix inversions. This cost could be significantly reduced by using the result from the previous iteration to warm-start a Newton-style iterative inversion routine. This approach was used by Anil et al. [2] to reduce the cost of computing inverse matrix roots as part of a wider deep learning optimisation scheme. We did not make use of such optimisations in our experiments.
Appendix H Experiment details
@@
The $x$-axis in Figure 2 of Section 4 corresponds to the number of NUTS steps, or more specifically, the number of NUTS candidate evaluations. This number is hardware and implementation agnostic, and roughly corresponds to the total computational cost incurred up to that point. Wall-clock time results are given in Appendix J.
@@
For EP, the step size ($\alpha$) was drawn log-uniformly in the range $(10^{-4},1)$. $n_{\text{samp}}$ was drawn log-uniformly between $[d+2.5,10000.5)$ and then rounded to the nearest integer, where $d$ is the dimensionality of $z$; note that the debiasing estimator used by Xu et al. [47] for MVN families is not defined for $n_{\text{samp}}\leq d+2$. The thinning ratio was drawn uniformly from $\{1,2,3,4\}$. For EP-$\eta$ and EP-$\mu$, the step size ($\epsilon$) was drawn log-uniformly in the range $(10^{-5},10^{-2})$, and $n_{\text{samp}}$ was fixed to $1$. For SNEP, the step size was drawn log-uniformly in the range $(10^{-5},10^{-2})$. $n_{\text{samp}}$ and $n_{\text{inner}}$ were (independently) drawn log-uniformly in the range [.$[.5,10.5)$ and then rounded to the nearest integer.
@@
All the methods considered can be used in a double-loop manner by taking $n_{\text{inner}}>1$ inner steps per outer update. We did not find this necessary in any of our experiments, and so we fixed $n_{\text{inner}}=1$ for EP, EP-$\eta$ and EP-$\mu$. For SNEP we included $n_{\text{inner}}$ in the search space (as described above) to make sure that setting it to $1$ was not negatively impacting its performance.
@@
Implementations were written in JAX [5], with NUTS [24] used as the underlying sampler. We used the numpyro [38] implementation of NUTS with default settings. For experiments with a NIW base family we performed mean-to-natural parameter conversions using the method of So [41], with JAXopt [4] used to perform implicit differentiation through the iterative solve.
Sampling hyperparameter adaptation
We performed regular warmup phases in order to adapt the samplers to the constantly evolving tilted distributions, consistent with prior work [47, 44]. The frequency and duration of these warmup phases was also included in the hyperparameter search as follows. We drew the duration (number of samples) of each warmup phase log-uniformly in the range $[99.5,1000.5)$ and then rounded to the nearest integer. We then drew the sampling-to-warmup ratio log-uniformly in the range $(1,4)$. This ratio then determined how frequently the warmup was performed, which we rounded to the nearest positive integer number of updates. Note that the Pareto frontier plots in Figures 2 and 6 include the computation/time spent during warmup phases.
@@
For EP, EP-$\eta$ and EP-$\mu$, we initialised the site parameters to $\lambda_{i}=\mathbf{0}\ \forall\ i$ for all models, with the exception of the cosmic radiation model for reasons we discuss later. SNEP is not compatible with improper site potentials, and so for models where $\mathcal{F}$ was MVN we used $\lambda_{i}=(2m)^{-1}\theta_{0}\ \forall\ i$, consistent with Vehtari et al. [44]. Unfortunately this initialisation strategy for SNEP is not always valid when $\mathcal{F}$ is NIW. This is because the site potentials in SNEP are restricted to being proper distributions, and this constraint was violated when using this initialisation in our experiments. We tried various other initialisation strategies, but none produced satisfactory results or allowed us to make a meaningful comparison, hence we omitted SNEP from the experiments with NIW $\mathcal{F}$; namely, the political survey hierarchical logistic regression and neural response model experiments.
@@
H.1 Hierarchical logistic regression
In these experiments, there were $m$ logistic regression problems, each with $n$ covariate/response pairs, so that $\mathcal{D}_{i}=\{(x_{i,j},y_{i,j})\}$, $x_{i,j}\in\mathbb{R}^{d}$, and $y_{i,j}\in\{0,1\}$. Each regression problem had its own (unobserved) vector of regression coefficients $w_{i}\in\mathbb{R}^{d}$, with prior density $p(w_{i}\mid z)$ parameterised by global parameters $z$.
@@ H.1 Hierarchical logistic regression
In the political survey data experiment, $\mathcal{F}$ was the family of normal-inverse-Wishart (NIW) distributions, and $z=(\mu,\text{vech}(\Sigma))$ was used to parameterise the group-level coefficient prior with density $p(w_{i}\mid z)=\mathcal{N}(w_{i}\mid\mu,\Sigma)$. $p_{0}$ was a NIW prior on $\mu$ and $\Sigma$, with parameters $\mu_{0}=\mathbf{0}$, $\nu=9$, $\lambda=1$, and $\Psi=I$ (following Wikipedia notation). We used a log-Cholesky parameterisation of $\Sigma$ for the purposes of sampling. The dataset consisted of binary responses to the statement "Allow employers to decline coverage of abortions in insurance plans (Support / Oppose)". We constructed $m=50$ regression problems, corresponding to the $50$ US states, and truncated the data so that there were exactly $n=97$ responses for each state. We used 6 predictor variables from the dataset, corresponding to characteristics of a given survey participant. The predictors were binary variables conveying: age (3 groups), ethicity (2 groups), education (3 groups) and gender. We also included a state-level intercept to capture variation in the base response level between states, so that $d=7$. This setup is based on one used by Lopez-Martin et al. [29]. The data is available at https://github.com/JuanLopezMartin/MRPCaseStudy.
H.2 Cosmic radiation model
@@
In this experiment, there were $m$ nonlinear regression problems, each corresponding to a model of the relationship between diffuse galactic far ultraviolet radiation (FUV) and 100-µm infrared (i100) emission in a particular section of the observable universe.
@@
Each nonlinear regression model was parameterised by parameters $w_{i}\in\mathbb{R}^{9}$. The $m$ regression problems were related through common hyperparameters $z\in\mathbb{R}^{18}$, which parameterised the section-level densities $p(\mathcal{D}_{i},w_{i}\mid z)$ where $\mathcal{D}_{i}$ is the observed data for sector $i$. $\mathcal{F}$ was the family of MVN distributions, and the prior on $z\in\mathbb{R}^{18}$ had density $p_{0}(z)=\mathcal{N}(\mathbf{0},10I)$. The specifics of $p(\mathcal{D}_{i},w_{i}\mid z)$ are quite involved and we refer the reader to Vehtari et al. [44] for further details.
@@
Vehtari et al. [44] applied this model to data obtained from the Galaxy Evolution Explorer telescope. We were unable to obtain the dataset, and so we generated synthetic data using hyperparameters parameters that were tuned by hand to try and match the qualitative properties of the original data set (see Appendix K for examples). We used a reduced number of $m=36$ sites and $n=200$ observations per site to reduce computation, allowing us to perform a comprehensive hyperparameter search.
@@
Using the conventional site parameter initialisation of $\lambda_{i}=\mathbf{0}\ \forall i$ resulted in most runs of EP failing during early iterations. We found that this was resolved by initialising with the method used by Vehtari et al. [44] for SNEP, that is, $\lambda_{i}=(2m)^{-1}\theta_{0}\ \forall i$, and so we used this initialisation for all methods. We note, however, that the performances of EP-$\eta$ and EP-$\mu$ were largely unaffected by this change.
H.3 Neural response model
@@
In this experiment $\mathcal{F}$ was the family of NIW distributions. $z=\bm{(}\mu,\text{vech}(\Sigma)\bm{)}\in\mathbb{R}^{65}$ was used to parameterise $\mathcal{N}(\log r_{j};\mu,\Sigma)$, the prior distribution density for latent log firing rates in trial $j$, for $j=1,\ldots,144$. $p_{0}$ was NIW with parameters $\mu_{0}=\mathbf{1}$, $\nu=12$, $\lambda=2.5$, and $\Psi=1.25I$ (again following Wikipedia notation). The observed spike count for neuron $k$ in trial $j$, $x_{j,k}\in\mathbb{N}$, was modelled as $\text{Poisson}\bm{(}c_{j,k},\exp(r_{j,k})\bm{)}$, for $j=1,\ldots,144$, $k=1,\ldots,10$. We again used a log-Cholesky parameterisation of $\Sigma$ for sampling.
We grouped trials together into $m=8$ batches of $n=18$ trials, so that $w_{i}=(\log r_{(i-1)180+1},\ldots,\log r_{(i-1)180+180})\in\mathbb{R}^{180}$ was the concatenation of (log) firing rates for all $10$ neurons across $18$ trials, with $\mathcal{D}_{i}=\{(x_{j,1},\ldots,x_{j,10})\}_{j=(i-1)180+1}^{(i-1)180+180}$ the observed spike counts.
Appendix I Hyperparameter sensitivity
In this appendix, we examine the effect of varying hyperparameters of EP and EP-$\eta$ on the synthetic hierarchical logistic regression experiment of Section 4.1. The results for EP-$\mu$ are similar to those of EP-$\eta$ and so we do not consider it separately here.
@@ Appendix I Hyperparameter sensitivity
In Figure 4a we show the effect of varying the number of samples used for estimating updates in EP; more accurate regions of the frontier generally require more samples. Figure 4b similarly shows the effect of varying the step size; more accurate regions of the frontier generally correspond to a smaller step size. Finally, Figure 4c shows the effect of varying the thinning ratio $\tau$; more accurate regions of the frontier correspond to more thinning. Together these plots illustrate the difficulty of tuning EP in stochastic settings. Tracing out the frontier is a three-dimensional problem. To achieve the best accuracy within the compute budget used for this experiment, the practitioner would need to set $400<n_{\text{samp}}\leq 500$, $0.1<\alpha\leq 0.3$, and $\tau=2$; and any deviation from this would seemingly result in suboptimal accuracy.
(a) Number of samples
(b) Step size
(c) Thinning ratio
Figure 4: The effect of varying EP hyperparameters; partial Pareto frontiers showing the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis).
@@
In Figure 5a we show the effect of varying the number of samples used for estimating updates in EP-$\eta$; the frontier is traced out with $n_{\text{samp}}=1$. We note however that the difference between $n_{\text{samp}}=1$ and $n_{\text{samp}}=10$ is relatively small, and so it may be sensible to choose $n_{\text{samp}}>1$ to make efficient use of parallel hardware, or to minimise per-iteration overheads (see Appendix G). Figure 5b similarly shows the effect of varying the step size; with more accurate regions of the frontier corresponding to a smaller step size. This also suggests that in practice it may make sense to set $\epsilon$ relatively large at first and then gradually decay it to improve the accuracy. Finally, Figure 5c shows the effect of varying $n_{\text{inner}}$, the number of inner steps performed per outer update. $n_{\text{inner}}>1$ corresponds to using EP-$\eta$ in "double-loop" mode. We did not find it necessary to increase $n_{\text{inner}}$ above one to obtain convergence in our experiments, but Figure 5c demonstrates that doing so would have a relatively small impact on performance. Together these plots illustrate that hyperparameter tuning for EP-$\eta$ is relatively straightforward. The frontier can be largely be traced out by varying $\epsilon$, and is relatively insensitive to $n_{\text{samp}}$ and $n_{\text{inner}}$.
(a) Number of samples
(b) Step size
(c) Number of inner steps
@@
Figure 5: The effect of varying EP-$\eta$ hyperparameters; partial Pareto frontiers showing the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis).
Appendix J Wall-clock time results
@@
In Figure 6, we show Pareto frontiers, with the y-axis showing the lowest KL divergence achieved by any hyperparameter setting, and the x-axis shows the cumulative number of seconds elapsed; that is, the wall-clock time equivalent of Figure 2. These results are in broad agreement with those of Figure 2, which suggests the sampling cost does indeed dominate the computational overheads of EP-$\eta$ and EP-$\mu$ in these experiments. The wall-clock time results for EP-$\eta$ and EP-$\mu$ would likely be improved by using more than one sample per update, by making more efficient use of hardware resources and minimising per-iteration overheads.
@@
We note that these times are necessarily implementation and hardware dependent. We did not make particular efforts to optimise for wall-clock time. In Appendix G we discuss approaches for minimising per-iteration overheads; these would likely improve wall-clock time performance for all methods, but should disproportionately favour EP-$\eta$ and EP-$\mu$, due to their frequent updates and larger per-iteration overheads compared to EP and SNEP.
(a) Hierarchical logistic regression: MVN prior
(b) Hierarchical logistic regression: NIW prior
(c) Cosmic radiation model
(d) Neural response model

### v2

@@ Appendix I Computational cost analysis
EP draws $n_{\text{samp}}$ samples from each of the $m$ sites. It then performs $m$ backward mappings, to map from moments back to updated site parameters. The total cost is then $mn_{\text{samp}}c_{\text{samp}}+mc_{\text{bwd}}$
@@
EP-$\eta$ draws $n_{\text{samp}}$ samples from each of the $m$ sites. Each update also involves one forward mapping per iteration. In addition, each iteration also requires $m$ JVPs through the backward mapping $\nabla A^{*}(.)$. The primals of this JVP are the same for each site, so the linearisation only needs to be performed once per update, costing $\approx 2c_{\text{bwd}}$, but we have $m$ tangents. The overall cost of an EP-$\eta$ iteration is therefore $\approx mn_{\text{samp}}c_{\text{samp}}+c_{\text{fwd}}+(2+m)c_{\text{bwd}}$.
@@
EP-$\mu$ draws $n_{\text{samp}}$ samples from each of the $m$ sites. Each update involves one forward mapping, and $m$ backward mappings per iteration, but notably does not require any JVPs. The overall cost of an EP-$\mu$ iteration is therefore $mn_{\text{samp}}c_{\text{samp}}+c_{\text{fwd}}+mc_{\text{bwd}}$.
@@
SNEP draws $n_{\text{samp}}$ samples from each of the $m$ sites, and performs $m$ backward mappings per iteration. The total cost is then $mn_{\text{samp}}c_{\text{samp}}+mc_{\text{bwd}}$, which is the same as EP.
In this work we largely assume that $c_{\text{samp}}$ is the dominant cost. This is often the case when the sampling is performed using MCMC. For example, in the NUTS sampler with default numpyro settings – as used in our evaluation – each sample can involve evaluating up to 1024 gradients of the tilted distribution log density.
The other costs, $c_{\text{fwd}}$ and $c_{\text{bwd}}$, depend on $\mathcal{F}$. In the case of a MVN family with dense covariance, the foward and backward mappings are $O(d^{3})$, for $d$ the dimensionality of $z$. When $d$ is large, $c_{\text{fwd}}$ and $c_{\text{bwd}}$ become more significant, in which case the balance will shift in favour of using more samples per update (and taking larger steps). When $d$ is very large, however, it may be necessary to choose a diagonal covariance base family instead, in which case the mappings can be performed in $O(d)$ time. Also note that when $d$ increases, $c_{\text{samp}}$ will also typically increase at a rate faster than $d$, even with optimal tuning [8]. In our evaluation we used just a single sample per update for EP-$\eta$ and EP-$\mu$, and they significantly outperformed the baselines when measure in both NUTS steps and wall-clock-time.
We also note that the per-iteration cost of $c_{\text{fwd}}$ and $c_{\text{bwd}}$ could be reduced by exploiting the results of previous iterations. For example, in the case of a MVN with dense covariance, the cost of the forward and backward mappings are dominated by matrix inversions. This cost could be significantly reduced by using the result from the previous iteration to warm-start a Newton-style iterative inversion routine. This approach was used by Anil et al. [2] to reduce the cost of computing inverse matrix roots as part of a wider deep-learning optimisation scheme. We did not make use of such optimisations in our experiments.
Appendix J Evaluation details
@@
The $x$-axis in Figure 2 of Section 4 corresponds to the number of NUTS steps, or more specifically, the number of NUTS candidate evaluations. This number is hardware and implementation agnostic, and roughly corresponds to the total computational cost incurred up to that point. Wall-clock time results are given in Appendix L.
@@
The step size for EP, $\alpha$, was drawn log-uniformly in the range $(10^{-4},1)$, and $n_{\text{samp}}$ was drawn log-uniformly between $[d+2.5,10000.5)$ and then rounded to the nearest integer, where $d$ is the dimensionality of $z$. Note that the debiasing estimator used by Xu et al. [52] for MVN families is not defined for $\smash{n_{\text{samp}}\leq d+2}$. The thinning ratio was drawn uniformly from $\{1,2,3,4\}$. The step size for EP-$\eta$ and EP-$\mu$, $\epsilon$, was drawn log-uniformly in the range $(10^{-5},10^{-2})$, and $n_{\text{samp}}$ was fixed to $1$. The step size for SNEP was drawn log-uniformly in the range $(10^{-5},10^{-2})$, with $n_{\text{samp}}$ and $n_{\text{inner}}$ (independently) drawn log-uniformly in the range [.$[.5,10.5)$ and then rounded to the nearest integer.
@@
All the methods considered can be used in a double-loop manner by taking $n_{\text{inner}}>1$ inner steps per outer update. We did not find this necessary in any of our experiments, and so we fixed $n_{\text{inner}}=1$ for EP, EP-$\eta$ and EP-$\mu$. For SNEP we included $n_{\text{inner}}$ in the search space (as described above) to ensure that setting it to $1$ was not negatively impacting its performance.
@@
Implementations were written in JAX [7], with NUTS [27] used as the underlying sampler. We used the numpyro [42] implementation of NUTS with default settings. For experiments with a NIW base family, we performed mean-to-natural parameter conversions using the method of So [45], with JAXopt [5] used to perform implicit differentiation through the iterative solve.
MCMC hyperparameter adaptation
We performed regular warm-up phases in order to adapt the samplers to the constantly evolving tilted distributions, consistent with prior work [52, 48]. The frequency and duration of these warm-up phases was also included in the hyperparameter search as follows. We drew the duration (number of samples) of each warm-up phase log-uniformly in the range $[99.5,1000.5)$ and then rounded to the nearest integer. We then drew the sampling-to-warm-up ratio log-uniformly in the range $(1,4)$. This ratio then determined how frequently the warm-up was performed, which we rounded to the nearest positive integer number of updates. Note that the Pareto frontier plots in Figures 2 and 7 include the computation / time spent during warm-up phases.
@@
For EP, EP-$\eta$ and EP-$\mu$, we initialised the site parameters to $\lambda_{i}=\mathbf{0}\ \forall\ i$ for all models, with the exception of the cosmic radiation model for reasons we discuss later. SNEP is not compatible with improper site potentials, and so for models where $\mathcal{F}$ was MVN we used $\lambda_{i}=(2m)^{-1}\eta_{0}\ \forall\ i$, consistent with Vehtari et al. [48]. Unfortunately this initialisation strategy for SNEP is not always valid when $\mathcal{F}$ is NIW. This is because the site potentials in SNEP are restricted to being proper distributions, and this constraint was violated when using this initialisation in our experiments. We tried various other initialisation strategies, but none produced satisfactory results or allowed us to make a meaningful comparison, hence we omitted SNEP from the experiments with NIW $\mathcal{F}$; namely, the political survey hierarchical logistic regression and neural response model experiments.
@@
J.1 Hierarchical logistic regression
In the hierarchical logistic regression (HLR) experiments, there were $m$ groups (logistic regression problems) each with $n$ covariate/response pairs, so that $\mathcal{D}_{i}=\{(x_{i,j},y_{i,j})\}$, $x_{i,j}\in\mathbb{R}^{d}$, and $y_{i,j}\in\{0,1\}$, for $i=1$ to $m$, and $j=1$ to $n$. Each group had its own unobserved vector of regression coefficients $w_{i}\in\mathbb{R}^{d}$, with density $p(w_{i}\mid z)$ parameterised by global parameters $z$.
@@
In the political survey data experiment, $\mathcal{F}$ was a normal-inverse-Wishart (NIW) family of distributions, and $z=(\mu,\text{vech}(\Sigma))$ was used to parameterise the group-level coefficient density $p(w_{i}\mid z)=\mathcal{N}(w_{i}\mid\mu,\Sigma)$. $p_{0}$ was a NIW prior on $\mu$ and $\Sigma$, with parameters $\mu_{0}=\mathbf{0}$, $\nu=9$, $\lambda=1$, and $\Psi=I$ (following Wikipedia notation). We used a log-Cholesky parameterisation of $\Sigma$ for the purposes of sampling. The dataset consisted of binary responses to the statement “Allow employers to decline coverage of abortions in insurance plans (Support / Oppose)”. We constructed $m=50$ regression problems, corresponding to the $50$ US states, and truncated the data so that there were exactly $n=97$ responses for each state. We used 6 predictor variables from the dataset, corresponding to characteristics of a given survey participant. The predictors were binary variables conveying: age (3 groups), ethicity (2 groups), education (3 groups) and gender. We also included a state-level intercept to capture variation in the base response level between states, so that $d=7$. This setup is based on one used by Lopez-Martin et al. [33]. The data is available at https://github.com/JuanLopezMartin/MRPCaseStudy.
J.2 Cosmic radiation model
@@
In this experiment, there were $m$ nonlinear regression problems, each corresponding to a model of the relationship between diffuse galactic far ultraviolet radiation (FUV) and 100-µm infrared (i100) emission in a particular sector of the observable universe.
@@
The nonlinear regression model for sector $i$ had parameters $w_{i}\in\mathbb{R}^{9}$. The $m$ regression problems were related through common hyperparameters $z\in\mathbb{R}^{18}$, which parameterised the section-level densities $p(\mathcal{D}_{i},w_{i}\mid z)$ where $\mathcal{D}_{i}$ is the observed data for sector $i$. $\mathcal{F}$ was the family of MVN distributions, and the prior on $z\in\mathbb{R}^{18}$ had density $p_{0}(z)=\mathcal{N}(\mathbf{0},10I)$. The specifics of $p(\mathcal{D}_{i},w_{i}\mid z)$ are quite involved and we refer the reader to Vehtari et al. [48] for further details.
@@
Vehtari et al. [48] applied this model to data obtained from the Galaxy Evolution Explorer telescope. We were unable to obtain the dataset, and so we generated synthetic data using hyperparameters that were tuned by hand to try and match the qualitative properties of the original data set (see Appendix N for examples). We used a reduced number of $m=36$ sites and $n=200$ observations per site to reduce computation, allowing us to perform a comprehensive hyperparameter search.
@@
Using the conventional site parameter initialisation of $\lambda_{i}=\mathbf{0}\;\forall\;i$ resulted in most runs of EP failing during early iterations. We found that this was resolved by initialising with the method used by Vehtari et al. [48] for SNEP, that is, $\lambda_{i}=(2m)^{-1}\eta_{0}\;\forall\;i$, and so we used this initialisation for all methods. We note, however, that the performances of EP-$\eta$ and EP-$\mu$ were largely unaffected by this change.
J.3 Neural response model
@@
$z=\bm{(}\mu,\text{vech}(\Sigma)\bm{)}\in\mathbb{R}^{65}$ was used to parameterise $\mathcal{N}(\log r_{j};\mu,\Sigma)$, the density for latent log firing rates in trial $j$, for $j=1,\ldots,144$. $p_{0}\in\mathcal{F}$ was NIW with parameters $\mu_{0}=\mathbf{1}$, $\nu=12$, $\lambda=2.5$, and $\Psi=1.25I$ (again following Wikipedia notation). The observed spike count for neuron $k$ in trial $j$, $x_{j,k}\in\mathbb{N}$, was modelled as $\text{Poisson}\bm{(}c_{j,k},\exp(r_{j,k})\bm{)}$, for $j=1,\ldots,144$, $k=1,\ldots,10$. We again used a log-Cholesky parameterisation of $\Sigma$ for sampling.
We grouped trials together into $m=8$ batches of $n=18$ trials, so that $\smash{w_{i}=(\log r_{180(i-1)+1},\ldots,\log r_{180i})\in\mathbb{R}^{180}}$ was the concatenation of log firing rates for all $10$ neurons across $18$ trials, with $\mathcal{D}_{i}=\{(x_{j,1},\ldots,x_{j,10})\}$ the observed spike counts, for $j=180(i-1)+1$ to $180i$.
Appendix K Hyperparameter sensitivity
In this appendix, we examine the effect of varying hyperparameters of EP and EP-$\eta$ on the synthetic hierarchical logistic regression experiment of Section 4. The results for EP-$\mu$ are similar to those of EP-$\eta$ and so we do not consider it separately here.
@@
In the left panel of Figure 5 we show the effect of varying the number of samples used for estimating updates in EP – more accurate regions of the frontier generally require more samples. The middle panel, similarly, shows the effect of varying the step size, and the right panel shows the effect of varying the thinning ratio $\tau$. Together these plots illustrate the difficulty of tuning EP in stochastic settings – tracing out the frontier is a three-dimensional problem. For example, to achieve the best accuracy within a compute budget of $10^{7}$ steps, the practitioner would need to set $400<n_{\text{samp}}\leq 500$, $0.1<\alpha\leq 0.3$, and $\tau=2$, and any deviation from this would seemingly result in suboptimal accuracy.
Figure 5: The effect of varying EP hyperparameters. Partial Pareto frontiers show the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis).
@@
In the left panel of Figure 6 we show the effect of varying the number of samples used for estimating updates in EP-$\eta$ – the frontier is traced out with $n_{\text{samp}}=1$. We note however that the difference between $n_{\text{samp}}=1$ and $n_{\text{samp}}=10$ is relatively small, and so it may be sensible to choose $n_{\text{samp}}>1$ to make efficient use of parallel hardware, or to minimise per-iteration overheads (see Appendix I). The middle panel, similarly, shows the effect of varying the step size, with more accurate regions of the frontier corresponding to a smaller step size. This also suggests that in practice it may make sense to set $\epsilon$ relatively large at first and then gradually decay it to improve the accuracy. Finally, the right panel shows the effect of varying $n_{\text{inner}}$, the number of inner steps performed per outer update. $n_{\text{inner}}>1$ corresponds to using EP-$\eta$ in “double-loop” mode. We did not find it necessary to increase $n_{\text{inner}}$ above one to obtain convergence in our experiments, but Figure 6 (right) demonstrates that doing so would have a relatively small impact on performance. Together these plots illustrate that hyperparameter tuning for EP-$\eta$ is relatively straightforward. The frontier can be largely be traced out by varying $\epsilon$, and is relatively insensitive to $n_{\text{samp}}$ and $n_{\text{inner}}$.
@@
Figure 6: The effect of varying EP-$\eta$ hyperparameters. Partial Pareto frontiers show the number of NUTS steps ($x$-axis) against the KL divergence from $p$ to an estimate of the optimum ($y$-axis).
Appendix L Pareto frontiers in wall-clock time
@@
In Figure 7, we show Pareto frontiers, with the y-axis showing the lowest KL divergence achieved by any hyperparameter setting, and the x-axis shows the cumulative number of seconds elapsed – that is, the wall-clock time equivalent of Figure 2. These results are in broad agreement with those of Figure 2, which suggests that the sampling cost does indeed dominate the computational overheads of EP-$\eta$ and EP-$\mu$ in these experiments. The wall-clock time results for EP-$\eta$ and EP-$\mu$ would likely be improved by using more than one sample per update, by making more efficient use of hardware resources and minimising per-iteration overheads.
@@
We note that these times are necessarily implementation and hardware dependent. We did not make particular efforts to optimise for wall-clock time. In Appendix I we discuss approaches for minimising per-iteration overheads. These would likely improve wall-clock time performance for all methods, but should disproportionately favour EP-$\eta$ and EP-$\mu$, due to their frequent updates and larger per-iteration overheads compared to EP and SNEP.